OpenStudy (anonymous):

Solving simultaneous differential equation? Solving simultaneous differential equation dx /dt = - xA / (x+y) dy /dt = -y B / ( x+ y ) A and B are constants x and y are functions of t . I want to know value of x and y in terms of constants and t . Please help and suggest literature. What should be matrix if I have to solve by eigen value method , what should be appoach please help .

OpenStudy (anonymous):

elementary differential equations by Boyce DiPrima To solve this things, i'll use Cayley-Hamilton theorem I don't know if you use numeric methods to solve this stuff or not so...Read xD

OpenStudy (anonymous):

But how to put into matrix form . As for following diff eqn form dx/dt = A x + B y dy / dt = Cx + D y matrix becomes A B C D eigenvalues and eigen vectors are calculated . What should be matrix in my my case as I have diff eqn dx /dt = - xA / (x+y) dy /dt = -y B / ( x+ y ) Please put light .

OpenStudy (dumbcow):

i don't think you can put it in matrix form because it is Non-linear your example is system of linear diff equations....the given system is non-linear so other methods must be used -I don't think it has an explicit solution for x(t) and y(t) you can make the inference that: \[\frac{dy}{dx} = \frac{dy}{dt}*\frac{dt}{dx}\] \[\frac{dy}{dx} = \frac{By}{Ax}\] thus \[y = x^{B/A}\] then by substitution \[\frac{dx}{dt} = \frac{-Ax}{x+x^{B/A}}\] simplify \[\frac{dx}{dt} = \frac{-A}{1+x^{B/A -1}}\] thus \[x + \frac{A}{B}x^{B/A} = -At\] (ignoring constants of integration) from here you are stuck isolating x(t) ? hope this helps here is wolfram solution http://www.wolframalpha.com/input/?i=x%27%28t%29+%3D+-Ax%2F%28x%2By%29+%2C+y%27%28t%29+%3D+-By%2F%28x%2By%29

OpenStudy (anonymous):

Thanks both Dumbcow and Umangiasd.

OpenStudy (anonymous):

In the step \[ \frac{ dy }{ dx } = \frac{ By }{ Ax } ,y = x^{\frac{ b }{ a }} \] or \[y = k(t)x ^{\frac{ b }{ a }}\] , k is substituted as a constant having logarithmic form . Will there be integration constant . I think that constant may be a function of t . Am I right or not .

OpenStudy (dumbcow):

@ankushsci yes you are right, i ignored the integration constant there, however it is just a constant NOT a function of t

OpenStudy (anonymous):

Thanks dumbcow.

OpenStudy (dumbcow):

yw